A mass of 3 kg is to rest vertically over a spring with a spring constant, k, of 1400 N/m. What is the energy absorbed by the spring?
1
Expert's answer
2019-07-19T14:14:44-0400

"k=\frac{F}{m}"

"x=\frac{F}{k}"

"x=\frac{mg}{k}"

"mg=3\times 9.81=29.43N"

"x=\frac{29.43}{1400}"

"x=0.021m"

Energy stored

"E=energy stored"

"E=0.5k\times x^2", "E=0.5\times 1400\times 0.021^2" "E=0.30 joule"

"Energy stored"